# The Absolute Anabelian Geometry of Virtual Curves Arising from Sections of Arithmetic Fundamental Groups of Configuration Spaces

### v1

@@ 6. Automorphisms of 2-Configuration Spaces
[cf. Definition 6.9(i)] are naturally isomorphic.
@@
Corollary 6.11 follows immediately from Proposition 5.3(ii) and the portion of Lemma 6.10(ii)(iii) concerning $[v,s]$. ∎
@@
In the situation of Proposition 6.1, suppose that $X$ is of type $(g,r)$. Then:
@@
(ii) Assume that $r\geq 1$ and $2g-2+r\geq 2$. Then for every $\alpha\in\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s])\setminus\{\text{\rm diag}_{X}\}$, after possibly passing to a finite field extension of $k$, there exists an open immersion $f:X\to Y$ obtained by removing a single element $\in Y(k)$ such that $\alpha$ is determined by $f$ (where we regard $\alpha$ as an element of $\text{\rm Cusp}(X)$ via the bijection of (1)). Moreover, $Y$ is of type $(g,r-1)$. Write $s^{\prime}\coloneqq f_{*}\circ s:G_{k}\to\Pi_{Y}$. Then the exact sequences
@@
[cf. Definition 6.9(i)] are naturally isomorphic.
@@
Assertion (i) follows immediately from the final portion of Proposition 6.1. Next, we consider assertion (ii). Observe that $\alpha$ determines an open immersion $f:X\to Y$ such that $f$, together with $s$, $s^{\prime}$, satisfy the conditions of Lemma 6.10. Therefore, assertion (ii) follows immediately from Corollary 6.11.
@@
$\mu\coloneqq\text{\rm Hom}(H^{2}(\Delta_{\mathcal{C}}^{\emptyset},\hat{\mathbb{Z}}),\hat{\mathbb{Z}})$
is a $\hat{\mathbb{Z}}$-module with $G_{k}$-action whose underlying topological group is isomorphic to $\hat{\mathbb{Z}}$. Moreover, there exists an open subgroup $H$ of $G_{k}$ such that $H$ acts on $\mu$ via the cyclotomic character.
@@
Without loss of generality, we may replace $k$ by a finite field extension of $k$. In particular, we may assume that $Y$ admits a $k$-rational point $y$. By Corollary 3.7 and Theorem 4.20(ii)(iii), $\Delta_{\mathcal{C}}^{\emptyset}$ may be naturally identified with $\Delta_{X_{y}}$. Therefore, Lemma 6.13 follows immediately from [[14], Proposition 1.2(ii)].
@@
In the situation of Proposition 6.1 [cf. also Proposition 5.2], assume further that $k$ is either a number field or a mixed-characteristic local field, and that $X$ is of type $(0,r)$ where $r\geq 3$.
@@
Since $\text{\rm diag}_{X}$ is fixed by the conjugation action of $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$, assertion (i) is equivalent to the following claim [cf. Theorem 4.20(ii), (iii)]:
@@
To verify assertion (ii), we consider the family of quotient groups $\{\Pi_{[\text{\rm pr}^{2/1}_{X},s]}^{S}\}$ of $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$, where $S$ ranges over the elements of $\text{\rm Cusp}^{3}([\text{\rm pr}^{2/1}_{X},s])$ [cf. Definition 6.9(ii)]. For any distinct elements $S,S^{\prime}\in\text{\rm Cusp}^{3}([\text{\rm pr}^{2/1}_{X},s])$ such that $\text{\rm Card}(S\cap S^{\prime})=2$, we claim that there exists a unique (up to an inner automorphism determined by an element of $\Delta_{[\text{\rm pr}^{2/1}_{X},s]}^{S}$) isomorphism $\theta^{S}_{S^{\prime}}:\Pi_{[\text{\rm pr}^{2/1}_{X},s]}^{S}\xrightarrow{\sim}\Pi_{[\text{\rm pr}^{2/1}_{X},s]}^{S^{\prime}}$ that satisfies the following condition:
(**∗ ∗): For any $\alpha\in\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s])$, $\theta^{S}_{S^{\prime}}$ maps $H^{S}_{[\text{\rm pr}^{2/1}_{X},s]}(\alpha)$ [cf. Definition 6.9(iii)] isomorphically onto $H^{S^{\prime}}_{[\text{\rm pr}^{2/1}_{X},s]}(\sigma(\alpha))$, where $\sigma$ is the unique element in $\text{\rm Aut}_{\text{\rm set}}(\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s]))$ that $\sigma$ permutes the two elements in $S\cap S^{\prime}$, permutes the two elements in $(S\cup S^{\prime})\setminus(S\cap S^{\prime})$, and fixes the other elements. Moreover, $\theta^{S}_{S^{\prime}}$ is compatible with the natural surjections $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}^{S}\twoheadrightarrow G_{k}$, $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}^{S^{\prime}}\twoheadrightarrow G_{k}$.
@@
To verify case (a), we note that, by Proposition 6.1, both $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}^{S}$ and $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}^{S^{\prime}}$ may be regarded as quotients (by some normal subgroup generated by elements in conjugacy classes of cuspidal inertia groups) of $\Pi_{[\text{\rm pr}^{1}_{X},s]}$. Therefore, case (a) follows immediately from Proposition 6.1 and Lemma 6.4(vi).
To verify case (b), by Corollary 6.11, we may assume that $X$ is of type $(0,3)$. Therefore, case (b) follows immediately from Proposition 6.8(vi). This completes the proof of existence. Now assertion (ii) follows immediately from Proposition 6.2, by taking the graph in the statement of assertion (ii) to be the graph whose vertices are the elements of $\text{\rm Cusp}^{3}([\text{\rm pr}^{2/1}_{X},s])$, and whose edges are the pairs $(S,S^{\prime})$ of distinct elements $S,S^{\prime}\in\text{\rm Cusp}^{3}([\text{\rm pr}^{2/1}_{X},s])$ such that $\text{\rm Card}(S\cap S^{\prime})=2$.
Assertion (iii) follows immediately from assertion (ii) and Proposition 6.2. Here, we also apply the compatibility stated in the final portion of (**∗ ∗), together with Lemma 6.13 and the “functoriality/compatibility” with respect to finite étale coverings discussed in [[16], Remark 1.10.1, (i)(ii)], to construct a cyclotome “$\mu$” as in Lemma 6.13 that is independent of the vertex $v$ and compatible with the isomorphisms $\theta^{S}_{S^{\prime}}$ of (**∗ ∗).
@@
Assertions (i), (ii), and (iii) follow immediately from Theorem 4.20(i)(ii), Proposition 5.2, and Theorem 5.13.
Next, we consider assertion (iv). Let $H$ (respectively, $H^{\prime}$) be the maximal open subgroup of $G_{k}$ (respectively, $G_{k^{\prime}}$) that acts trivially on $\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s])$ (respectively, $\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X^{\prime}},s^{\prime}])$). By (iii), $f$ induces an isomorphism between $H$ and $H^{\prime}$. Let $k_{H}$ (respectively, $k^{\prime}_{H^{\prime}}$) be the field determined by $H$ (respectively, $H^{\prime}$). By Theorem 6.14(iii), $f$ induces a natural isomorphism $k_{H}\xrightarrow{\sim}k^{\prime}_{H^{\prime}}$ that is compatible with the respective natural actions of $G_{k}/H$ and $G_{k^{\prime}}/H^{\prime}$ on $k_{H}$ and $k^{\prime}_{H^{\prime}}$, relative to the isomorphism $G_{k}/H\xrightarrow{\sim}G_{k^{\prime}}/H^{\prime}$ induced by $f$ [cf. assertion (i)]. By Galois theory, we have natural isomorphisms $k\xrightarrow{\sim}(k_{H})^{G_{k}/H}$, $k^{\prime}\xrightarrow{\sim}(k^{\prime}_{H^{\prime}})^{G_{k^{\prime}}/H^{\prime}}$. Therefore, $f$ induces a natural isomorphism $k\xrightarrow{\sim}k^{\prime}$ of fields.
@@
Case ($1$) follows immediately from Theorem 6.14(ii)(iii). To verify case ($2$), we begin by observing that it follows from Proposition 6.1 that both $X$ and $X^{\prime}$ admit a $k$-rational point, hence that $K(X)$ (respectively, $K(X^{\prime})$) is isomorphic over $k$ (respectively, $k^{\prime}$) to the function field of $\mathbf{P}^{1}_{k}$ (respectively, $\mathbf{P}^{1}_{k^{\prime}}$). Thus, case ($2$) follows immediately from assertion (iv). To verify case ($3$), we observe that $f$ induces an isomorphism $\Pi^{E}_{[\text{\rm pr}^{2/1}_{X},s]}\xrightarrow{\sim}\Pi^{E^{\prime}}_{[\text{\rm pr}^{2/1}_{X^{\prime}},s^{\prime}]}$, where we note that both the domain and the codomain of this isomorphism is a quotient by cuspidal inertia subgroups of the quotient of “$\Pi_{[\text{\rm pr}^{{2/1}}_{X},s]}$” by “$H$” constructed in Proposition 6.1. Therefore, case ($3$) follows immediately from Proposition 6.1 and Proposition 6.2.

### v2

@@ 6. Automorphisms of 2-Configuration Spaces
[cf. Definition˜6.9(i)] are naturally isomorphic.
@@
Corollary˜6.11 follows immediately from Proposition˜5.3(ii) and the portion of Lemma˜6.10(ii)(iii) concerning $[v,s]$. ∎
@@
In the situation of Proposition˜6.1, suppose that $X$ is of type $(g,r)$. Then:
@@
(ii) Assume that $r\geq 1$ and $2g-2+r\geq 2$. Then for every $\alpha\in\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s])\setminus\{\text{\rm diag}_{X}\}$, after possibly passing to a finite field extension of $k$, there exists an open immersion $f:X\to Y$ obtained by removing a single element $\in Y(k)$ such that $\alpha$ is determined by $f$ (where we regard $\alpha$ as an element of $\text{\rm Cusp}(X)$ via the bijection of (1)). Moreover, $Y$ is of type (g,r1)(g,r-1). Write $s^{\prime}\coloneqq f_{*}\circ s:G_{k}\to\Pi_{Y}$. Then the exact sequences
@@
[cf. Definition˜6.9(i)] are naturally isomorphic.
@@
Assertion (i) follows immediately from the final portion of Proposition˜6.1. Next, we consider assertion (ii). Observe that $\alpha$ determines an open immersion $f:X\to Y$ such that $f$, together with $s$, $s^{\prime}$, satisfy the conditions of Lemma˜6.10. Therefore, assertion (ii) follows immediately from Corollary˜6.11.
@@
$)\mu\coloneqq\text{\rm Hom}(H^{2}(\Delta_{\mathcal{C}}^{\emptyset},\hat{\mathbb{Z}}$$\hat{\mathbb{Z}})\$
is a $^\hat{\mathbb{Z}}$-module with $G_{k}$-action whose underlying topological group is isomorphic to $\hat{\mathbb{Z}}$. Moreover, there exists an open subgroup $H$ of $G_{k}$ such that $H$ acts on $\mu$ via the cyclotomic character.
@@
Without loss of generality, we may replace $k$ by a finite field extension of $k$. In particular, we may assume that $Y$ admits a $k$-rational point $y$. By Corollary 3.7 and Theorem 4.20(ii)(iii), $\Delta_{\mathcal{C}}^{\emptyset}$ may be naturally identified with $\Delta_{X_{y}}$. Therefore, Lemma˜6.13 follows immediately from [[13], Proposition 1.2(ii)].
@@
In the situation of Proposition˜6.1 [cf. also Proposition˜5.2], assume further that $k$ is either a number field or a mixed-characteristic local field, and that $X$ is of type $(0,r)$ where $r\geq 3$.
@@
Since $\text{\rm diag}_{X}$ is fixed by the conjugation action of $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$, assertion (i) is equivalent to the following claim [cf. Theorem˜4.20(ii), (iii)]:
@@
To verify assertion (ii), we consider the family of quotient groups $\{\Pi_{[\text{\rm pr}^{2/1}_{X},s]}^{S}\}$ of $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$, where $S$ ranges over the elements of $\text{\rm Cusp}^{3}([\text{\rm pr}^{2/1}_{X},s])$ [cf. Definition˜6.9(ii)]. For any distinct elements $S,S^{\prime}\in\text{\rm Cusp}^{3}([\text{\rm pr}^{2/1}_{X},s])$ such that $\text{\rm Card}(S\cap S^{\prime})=2$, we claim that there exists a unique (up to an inner automorphism determined by an element of $\Delta_{[\text{\rm pr}^{2/1}_{X},s]}^{S}$) isomorphism $\theta^{S}_{S^{\prime}}:\Pi_{[\text{\rm pr}^{2/1}_{X},s]}^{S}\xrightarrow{\sim}\Pi_{[\text{\rm pr}^{2/1}_{X},s]}^{S^{\prime}}$ that satisfies the following condition:
(**): For any $\alpha\in\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s])$, $\theta^{S}_{S^{\prime}}$ maps $H^{S}_{[\text{\rm pr}^{2/1}_{X},s]}(\alpha)$ [cf. Definition 6.9(iii)] isomorphically onto $H^{S^{\prime}}_{[\text{\rm pr}^{2/1}_{X},s]}(\sigma(\alpha))$, where $\sigma$ is the unique element in $\text{\rm Aut}_{\text{\rm set}}(\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s]))$ that $\sigma$ permutes the two elements in $S\cap S^{\prime}$, permutes the two elements in $(S\cup S^{\prime})\setminus(S\cap S^{\prime})$, and fixes the other elements. Moreover, $\theta^{S}_{S^{\prime}}$ is compatible with the natural surjections $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}^{S}\twoheadrightarrow G_{k}$, $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}^{S^{\prime}}\twoheadrightarrow G_{k}$.
@@
To verify case (a), we note that, by Proposition˜6.1, both $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}^{S}$ and $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}^{S^{\prime}}$ may be regarded as quotients (by some normal subgroup generated by elements in conjugacy classes of cuspidal inertia groups) of $\Pi_{[\text{\rm pr}^{1}_{X},s]}$. Therefore, case (a) follows immediately from Proposition˜6.1 and Lemma˜6.4(vi).
To verify case (b), by Corollary˜6.11, we may assume that $X$ is of type $(0,3)$. Therefore, case (b) follows immediately from Proposition˜6.8(vi). This completes the proof of existence. Now assertion (ii) follows immediately from Proposition 6.2, by taking the graph in the statement of assertion (ii) to be the graph whose vertices are the elements of $\text{\rm Cusp}^{3}([\text{\rm pr}^{2/1}_{X},s])$, and whose edges are the pairs $(S,S^{\prime})$ of distinct elements $S,S^{\prime}\in\text{\rm Cusp}^{3}([\text{\rm pr}^{2/1}_{X},s])$ such that $\text{\rm Card}(S\cap S^{\prime})=2$.
Assertion (iii) follows immediately from assertion (ii) and Proposition 6.2. Here, we also apply the compatibility stated in the final portion of (**), together with Lemma 6.13 and the “functoriality/compatibility” with respect to finite étale coverings discussed in [[16], Remark 1.10.1, (i)(ii)], to construct a cyclotome $\mu$” as in Lemma 6.13 that is independent of the vertex $v$ and compatible with the isomorphisms $\theta^{S}_{S^{\prime}}$ of (**).
@@
Assertions (i), (ii), and (iii) follow immediately from Theorem˜4.20(i)(ii), Proposition 5.2, and Theorem˜5.13.
Next, we consider assertion (iv). Let $H$ (respectively, $H^{\prime}$) be the maximal open subgroup of $G_{k}$ (respectively, $G_{k^{\prime}}$) that acts trivially on $\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s])$ (respectively, $\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X^{\prime}},s^{\prime}])$). By (iii), $f$ induces an isomorphism between $H$ and $H^{\prime}$. Let $k_{H}$ (respectively, $k^{\prime}_{H^{\prime}}$) be the field determined by $H$ (respectively, $H^{\prime}$). By Theorem˜6.14(iii), $f$ induces a natural isomorphism $k_{H}\xrightarrow{\sim}k^{\prime}_{H^{\prime}}$ that is compatible with the respective natural actions of $G_{k}/H$ and $G_{k^{\prime}}/H^{\prime}$ on $k_{H}$ and $k^{\prime}_{H^{\prime}}$, relative to the isomorphism $G_{k}/H\xrightarrow{\sim}G_{k^{\prime}}/H^{\prime}$ induced by $f$ [cf. assertion (i)]. By Galois theory, we have natural isomorphisms $k\xrightarrow{\sim}(k_{H})^{G_{k}/H}$, $k^{\prime}\xrightarrow{\sim}(k^{\prime}_{H^{\prime}})^{G_{k^{\prime}}/H^{\prime}}$. Therefore, $f$ induces a natural isomorphism $k\xrightarrow{\sim}k^{\prime}$ of fields.
@@
Case ($1$) follows immediately from Theorem˜6.14(ii)(iii). To verify case ($2$), we begin by observing that it follows from Proposition˜6.1 that both $X$ and $X^{\prime}$ admit a $k$-rational point, hence that $K(X)$ (respectively, $K(X^{\prime})$) is isomorphic over $k$ (respectively, $k^{\prime}$) to the function field of $\mathbf{P}^{1}_{k}$ (respectively, $\mathbf{P}^{1}_{k^{\prime}}$). Thus, case ($2$) follows immediately from assertion (iv). To verify case ($3$), we observe that $f$ induces an isomorphism $\Pi^{E}_{[\text{\rm pr}^{2/1}_{X},s]}\xrightarrow{\sim}\Pi^{E^{\prime}}_{[\text{\rm pr}^{2/1}_{X^{\prime}},s^{\prime}]}$, where we note that both the domain and the codomain of this isomorphism is a quotient by cuspidal inertia subgroups of the quotient of $\Pi_{[\text{\rm pr}^{{2/1}}_{X},s]}$” by “HH” constructed in Proposition˜6.1. Therefore, case ($3$) follows immediately from Proposition˜6.1 and Proposition˜6.2.